a capital of 500 dollars earns 150 dollars of interest in 6 years. What was the

interest rate if compound interest is used? What if simple interest is used?


1
Expert's answer
2021-06-15T18:23:16-0400

Compound interest

"A = P(1 +\frac{ r}{n})^{nt}"

A=final amount

P=initial principal balance

r=interest rate

n=number of times interest applied per time period

t=number of time periods elapsed

Amount=Interest+Amount

"=150+500=650"

"A = P(1 +\frac{ r}{n})^{nt}"

"650 = 500(1 +\frac{ r}{1})^{1×6}"

"1.3=(1+r)^6"

"\sqrt[6]{1.3}=1+r\\=1.0446750"

"1.0446750-1=0.0446750\\4.46750\%"

Simple interest

"I=P\u00d7I\u00d7T"

  • P = Principal Amount
  • I = Interest Amount
  • r = Rate of Interest per year in

"150=500×I×6\\150=3000I\\I=0.05\\0.05×100=5\%"